EXPLORING THE PHOTOREDUCTION OF Au(III) COMPLEXES IN THE GAS-PHASE

\underline{JESSE C. MARCUM}, SYDNEY H. KAUFMAN, J. MATHIAS WEBER, JILA, NIST, Department of Chemistry and Biochemistry, The University of Colorado, Boulder, CO.

We have used photodissociation spectroscopy to probe the electronic structure and photoreduction of Au(III) in gasphase complexes containing Cl^{−} and OH^{−}. The gas-phase electronic spectrum of [AuCl_{4}]^{−} closely resembles the aqueous solution spectrum, showing a lack of strong solvatochromic shifts. Substitution of Cl^{−} ligands with OH^{−} results in a strong blue shift, in agreement with ligand-field theory. Upon excitation, [AuCl_{4}]^{−} can dissociate by loss of either one or two neutral Cl atoms, resulting in the reduction of gold from Au(III) to Au(II) and Au(I) respectively. The hydroxide substituted complex, [AuCl_{2}(OH)_{2}]^{−}, demonstrates similar behavior but the only observable fragment channel is the loss of two neutral OH ligands, leading only to Au(I).